OpenStudy (anonymous):

Prove that all values of 〖(1-i) 〗^(i√2) lie on a straight line. Were i is the imaginary number √(-1).

OpenStudy (anonymous):

is that\[\large (1-i)^{i\sqrt{2}}\]?

OpenStudy (anonymous):

Yes

OpenStudy (experimentx):

upon simplification, we get \[ \huge \left( \sqrt{2} e^{i\pi\over 4}\right )^{i\sqrt2}\] using this probably you will find the principle root.

OpenStudy (experimentx):

do you need to find all the roots?

OpenStudy (anonymous):

Yes please

OpenStudy (experimentx):

I don't know right now ... let me ask some other people. most likely I'm thinking ... it has only one root.

OpenStudy (anonymous):

Thank you. I will post another problem in complex analysis.

OpenStudy (experimentx):

\[ \huge \left( e^{\ln \sqrt2+{i\pi\over 4}}\right )^{i\sqrt2} = e^{{-\sqrt 2 \over 4} + i\sqrt 2 \ln 2}\]

OpenStudy (anonymous):

Many Thanks

OpenStudy (experimentx):

no probs ...hardly i put any effort.